# REALM-Bench: A Real-World Planning Benchmark for LLMs and Multi-Agent Systems

### v1

@@ Abstract.
This benchmark suite provides a comprehensive evaluation framework for assessing both individual LLMs and multi-agent systems in real-world planning scenarios. The suite encompasses eleven designed problems that progress from basic to highly complex, incorporating key aspects such as multi-agent coordination, inter-agent dependencies, and dynamic environmental disruptions. Each problem can be scaled along three dimensions: the number of parallel planning threads, the complexity of inter-dependencies, and the frequency of unexpected disruptions requiring real-time adaptation. The benchmark includes detailed specifications, evaluation metrics, and baseline implementations using contemporary frameworks like LangGraph, enabling rigorous testing of both single-agent and multi-agent planning capabilities. Through standardized evaluation criteria and scalable complexity, this benchmark aims to drive progress in developing more robust and adaptable AI planning systems for real-world applications.
LLMs, Multi-Agent Systems, Planning Benchmark
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Conf; February; 2025isbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
As large language models (LLMs) continue to advance in reasoning and planning, as demonstrated by OpenAI’s GPT-4o-Task (OpenAI, 2024), DeepSeek’s R1 (DeepSeek-AI et al., 2025), Anthropic’s Claude 3.5 Sonnet (Anthropic, 2024), and Gemini (Team et al., 2023), the research community is increasingly focusing on developing multi-agent systems (MAS) powered by these models. Recent innovations include AutoGen (Wu et al., 2024), CAMEL (Li et al., 2023), CrewAI (Moura, 2024), LangGraph (LangChain AI, 2024), Dspy (Khattab et al., 2023), and XAgent (XAgent Team, 2023), among others. Although individual LLMs demonstrate significant capabilities, their true potential is realized when they collaborate to tackle complex real-world problems (Chang, 2024).
Most AI benchmarks emphasize perception, language understanding, or basic reasoning. However, real-world challenges, such as supply chain management, disaster response, healthcare logistics, and investment strategies, demand coordinated planning and decision-making among specialized agents. There is a pressing need for robust benchmarks that can evaluate the performance of both single-agent systems and MAS in these complex, high-stakes domains.
@@ 1.1. The REALM Benchmark Suite
REALM-Bench (Real-world Planning Benchmark for LLMs and Multi-Agent Systems) addresses the need for rigorous evaluation with carefully curated planning challenges. These scenarios are designed to be both tractable, enabling human validation and debugging, and sufficiently complex to push the boundaries of current AI systems. Each challenge requires reasoning and validation over sequential actions parallel processes, resource constraints, and unexpected disruptions (Chang, 2025, 2023b, 2023a; Chen et al., 2024; Huang et al., 2024).
The suite consists of eleven scenarios that progressively increase in complexity across three key dimensions:
@@
    Parallel Planning Threads: The number of concurrent planning processes that must be coordinated.
@@
    Inter-Dependencies: The complexity of relationships and constraints between these planning threads.
@@
    Disruption Frequency and Impact: The rate and severity of unexpected events that require the adaptation of the plan.
Next, we describe how each scenario can be scaled along these three dimensions.
1.2. Benchmark Scalability
While the base versions of each scenario enable detailed analysis and debugging, they can be scaled along the three dimensions defined above: parallel planning threads, inter-dependencies, and disruption frequency and impact.
For example, an urban ride-sharing scenario becomes increasingly complex as the number of vehicles and passengers grows, with interdependent carpooling routes and frequent traffic disruptions necessitating real-time plan adjustments.
This scalability allows AI planning systems to be evaluated under progressively challenging conditions, while still allowing detailed analysis of failure modes in simpler scenarios.
1.3. Availability and Access
The REALM-Bench Suite V1.0 is available on GitHub (Geng and Chang, 2025). In addition, we plan to host competitions and workshops at major AI conferences in 2025 to foster community engagement and further development.
2. Related Benchmark
Planning benchmarks have evolved from testing basic STRIPS-style planning to evaluating increasingly sophisticated planning capabilities. The International Planning Competition (IPC) has been a primary driver of planning benchmarks since 1998, using PDDL to specify domains like BlocksWorld, Logistics, and Rovers (Taitler et al., 2024). While valuable for testing classical planning algorithms, these benchmarks focus on deterministic environments with complete information and lack the dynamic disruptions common in real-world scenarios.
More recent benchmarks, such as the Process Planning Competition (PPC), have shifted toward continuous processes and temporal constraints (Smith and Johnson, 2020). Their manufacturing scenarios include parallel activities and resource dependencies, but the disruptions remain limited to machine breakdowns with known repair distributions. Similarly, the Dynamic Planning Competition introduces environmental changes during plan execution, yet it focuses primarily on path planning and navigation scenarios (Doe and Roe, 2022).
The annual Automated Negotiation Agents Competition (ANAC), established in 2010, has evolved to incorporate planning elements within its supply chain scenarios (Miller and Davis, 2023). However, its scope remains primarily focused on bilateral negotiations rather than comprehensive planning under uncertainty. For example, the 2024–25 competition featured a main challenge titled “Split the Pie,” an artificial yet simplified negotiation scenario where agents divide resources between parties. The supply chain problems in ANAC do not involve contingency planning, resource reallocation, or adaptation to unexpected disruptions.
Specifically for testing LLMs’ planning capabilities, TimeBench (Chu et al., 2023) and TaskBench (Shen et al., 2023) represent two approaches to evaluating AI planning. TimeBench focuses on temporal reasoning by testing systems’ ability to understand time dependencies and scheduling constraints, though it often relies on synthetic scenarios that fail to capture the dynamic nature of real-world temporal relationships, where deadlines shift and durations remain uncertain. TaskBench, on the other hand, evaluates practical task automation and step-by-step planning; it provides valuable insights into an AI system’s ability to decompose complex goals into manageable steps, but its scenarios may oversimplify the challenges of real-world automation, where outcomes are uncertain and processes are deeply interconnected.
This landscape reveals several gaps in existing benchmarks:
  1. 1.
    Limited Disruption Modeling: Most benchmarks treat uncertainties as static probability distributions rather than dynamic, interdependent events that can cascade through systems.
  2. 2.
    Simplified Dependencies: Real-world planning problems involve rich networks of temporal, resource, and causal dependencies that exceed the complexity found in current benchmarks.
  3. 3.
    Restricted Scope: Benchmarks tend to focus on specific subproblems (path planning, task allocation, etc.) rather than end-to-end planning scenarios that combine multiple challenges.
@@ 2. Related Benchmark
    Artificial Constraints: Many benchmarks use simplified representations (like PDDL) that cannot capture the nuanced constraints and objectives found in real-world planning problems.
@@
    Limited Scalability: Few benchmarks allow systematic scaling of complexity along multiple dimensions while maintaining problem tractability for analysis.
  6. 6.
    LLM Specific Challenges: Although LLMs have achieved remarkable successes, the transformer architecture exhibits certain limitations. For example, an attention sink phenomenon can cause certain tokens to be neglected, potentially skewing model predictions (Xiao et al., 2024). Additionally, maximum likelihood training can introduce biases that limit output diversity and quality (Holtzman et al., 2020; Chang, 2023b). Finally, chain-of-thought approaches may suffer from pitfalls such as error propagation and inconsistent reasoning (Brown et al., 2020; Prystawski et al., 2023). A test suite should specifically examine these LLM-related issues.
Our proposed REALM-Bench addresses these limitations by providing scenarios that combine dependencies, dynamic disruptions, and scalability while remaining tractable for systematic evaluation. This allows testing of planning systems under conditions that better reflect the challenges of real-world applications.
3. Benchmark Structure
Problems are categorized into three difficulty levels based on the number of parallel execution threads, the complexity of dependencies, and real-time disruptions.
3.1. Entry Level (1-2 threads)
Problems focusing on basic coordination with limited dependencies:
    Single or dual thread execution
    Basic timing and resource constraints
    Simple disruption scenarios
    Example: Campus tour coordination with one and two groups
3.2. Intermediate (3-4 threads)
Problems requiring significant coordination across multiple execution paths:
    Three to four parallel threads
    Complex timing dependencies
    Resource sharing constraints
    Example: Wedding logistics with multiple vehicles and tasks
3.3. Advanced (5+ threads)
Problems with real-world complexity:
    Five or more parallel threads
    Complex inter-dependencies
    Multiple resource conflicts
    Dynamic disruption scenarios
    Example: Thanksgiving dinner coordination, natural disaster relief, and supply chain management
3.4. Evaluation Metrics
Each problem is evaluated across five key dimensions:
    Planning Quality: Effectiveness of initial plan generation
    Coordination: Management of parallel thread execution
    Adaptation: Response to disruptions and changes
    Resource Management: Resolution of resource conflicts
    Constraint Satisfaction: Maintenance of problem constraints
@@ 4. Benchmark Problem Specifications
Metrics: - Total tour time: Minimize while meeting all constraints - Visit coverage: All locations must be visited Locations: Five campus buildings $L=\{S,L,B,A,D\}$ $S$: Student Center - $L$: Library - $B$: Lab Building - $A$: Athletics Center - $D$: Dormitory Travel Times: (minutes) $S\text{-}L:10,S\text{-}B:15,S\text{-}A:20,S\text{-}D:15$ $L\text{-}B:10,L\text{-}A:25,L\text{-}D:20$ $B\text{-}A:15,B\text{-}D:25$ $A\text{-}D:20$ Visit Requirements: - Start time given - Each location requires 30-minute visit - Tour starts/ends at Student Center ($S$) - Group size: 20 people Time Constraints: - Lab Building: Only 9 AM - 4 PM - Library: After 10 AM - Total tour must complete by 5 PM
@@
Groups: - $G_{1}$: 15 people (domestic students) - $G_{2}$: 20 people (international students) Locations: Ten campus buildings $L=\{S,L,B,A,D,C,M,R,H,P\}$ with capacities $cap_{l}$: $cap_{l}=\begin{cases}40&l\in\{S,A\}\\ 30&l\in\{L,D,C\}\\ 25&l\in\{B,M\}\\ 20&l\in\{R,H,P\}\end{cases}$ Constraints: - Tour start time for each group: between 9 AM and 10 AM - Total visitors must not exceed location capacity - Each location requires 30-minute visit - Both tours start/end at Student Center - Complete all tours by 5 PM Additional Requirements: - Lab tours ($B$) only 9 AM - 4 PM - Dining ($C$) must be visited between 11 AM - 2 PM - Library ($L$) after 10 AM
REALM-Bench comprises eleven foundational problem frameworks that systematically evaluate both sequential and reactive planning. Building on the key dimensions introduced earlier (parallel threads, inter-dependencies, and dynamic disruptions), these frameworks progress from straightforward single-thread execution to complex multi-agent scenarios with real-time challenges.
Considerations #1 Problem Complexity:
Each framework can be further scaled to create more challenging variants:
  • @itemi
    Expanding the scale of agents and resources (e.g., from dozens to thousands)
  • @itemi
    Widening geographic distribution (e.g., local to global)
  • @itemi
    Increasing disruption frequency and severity (e.g., isolated events to cascading failures)
  • @itemi
    Introducing uncertainties in execution times and outcomes (e.g., probabilistic durations)
  • @itemi
    Adding hierarchical dependency networks (e.g., sub-networks with internal dependencies)
  • @itemi
    Accounting for agent properties (e.g., atomicity, idempotency)
This extensible design enables researchers to incrementally assess their systems’ capabilities while preserving each scenario’s fundamental planning challenge.
Considerations #2 Implementation:
Three main approaches can be used to tackle these problems:
  • Rudimentary Manual Approach: Hand-craft a directed graph, choose a solver, and input parameter values to obtain a solution. Although direct, this method demands considerable human effort and domain expertise.
  • LLM-Assisted Semi-automation: Use an LLM to suggest algorithms, extract parameters from specifications, and assist with coding. This approach can handle moderate complexity but becomes less feasible at scale.
  • Full Automation: Provide the problem statement to a multi-agent framework (e.g., MACI (Chang, 2025)) that executes end-to-end, including validation and replanning. This approach, combined with human oversight, scales best for complex scenarios with large numbers of nodes and frequent environmental changes.
@@
Metrics: - On-time performance: No penalty for early arrivals. - Total distance traveled. Locations: Seven locations: $V=\{A,B,C,D,E,F,G\}$, where $G$ is Boston Logan Airport (BOS). Urban locations $A$$F$ are all 10 km of each other, while distances to BOS are 30+ km. $\begin{bmatrix}&A&B&C&D&E&F\\ A-F&10&10&10&10&10&10\\ \rightarrow G&35&33&36&34&32&31\end{bmatrix}$ Travel speed: ($A$$F$) 60 km/h, and ($A$$F\rightarrow G$) 100 km/h. Passengers: Each passenger specifies an arrival time at BOS ($G$). The dispatcher will instruct drivers when to pick up passengers to ensure on-time arrival at BOS. Ride Requests (Desired BOS arrival time given): - $r_{1}$: Pickup at $A$, to $G$ by 08:45     - $r_{2}$: Pickup at $B$, to $G$ by 08:50 - $r_{3}$: Pickup at $C$, to $G$ by 08:55     - $r_{4}$: Pickup at $D$, to $G$ by 09:00 Available Vehicles (Capacity 2 passengers): - $k_{1}$: at $A$, $k_{2}$: at $C$, and $k_{3}$: at $E$ Scheduling Constraints: - The dispatcher determines the pickup times based on a feasible schedule. Pickup times must allow the driver to first reach the passenger location ($A$ $F$) and then drive to $G$ in time.
4.1. P1: Campus Single-Tour Navigation
Problem Statement: A single autonomous agent must navigate a predefined campus environment to complete a sequence of waypoints while minimizing travel time. The scenario assumes a static environment without disruptions.

### v2

# REALM-Bench: A Benchmark for Evaluating Multi-Agent Systems on Real-world, Dynamic Planning and Scheduling Tasks

@@ Abstract.
This benchmark suite provides a comprehensive evaluation framework for assessing both individual LLMs and multi-agent systems in Real-world planning and scheduling scenarios. The suite encompasses 14 designed planning and scheduling problems that progress from basic to highly complex, incorporating key aspects such as multi-agent coordination, inter-agent dependencies, and dynamic environmental disruptions. Each problem can be scaled along three dimensions: the number of parallel planning threads, the complexity of inter-dependencies, and the frequency of unexpected disruptions requiring Real-time adaptation. The benchmark includes 14 detailed problem specifications, 15 comparison methods including Random, LPT, SPT, STPT, MPSR, DRL-Liu, GP, GEP, LSO, SPT/TWKR, DRL-Chen, DRL-Zhang, 2+ evaluation metrics, and baseline implementations using 3+ LLMs including GPT-4o, Claude-3.7, DeepSeek-R1, and 4 contemporary frameworks including LangGraph, AutoGen, CrewAI, and Swarm, enabling rigorous testing of both single-agent and multi-agent planning capabilities. Through standardized evaluation criteria and scalable complexity, this benchmark aims to be opened to public, and drive progress in developing more adaptable, robust, and scalable AI planning systems for Real-world applications.
Code:https://github.com/genglongling/REALM-Bench
Datasets:https://github.com/genglongling/M-APPLE-OS
LLMs, Multi-Agent Systems, Planning and Scheduling
conference: 2025; under review;
@@ 1. Introduction
Large language models (LLMs) today excel at single-turn interactions: answering questions, generating text, solving isolated reasoning problems, and providing explanations within individual conversations. However, real-world applications demand capabilities that current LLMs struggle with: multi-step planning over extended horizons, coordinated multi-agent collaboration, adaptive responses to dynamic environments, and sustained problem-solving that requires persistent state management and iterative refinement.
Recent advances in LLMs, including OpenAI’s GPT-4o-Task (OpenAI, 2024), DeepSeek’s R1 (DeepSeek-AI et al., 2025), Anthropic’s Claude 3.5 Sonnet (Anthropic, 2024), and Gemini (Team et al., 2023), demonstrate impressive natural language capabilities but remain limited when faced with complex, multi-faceted problems that require orchestrating multiple agents, maintaining coherent plans across disruptions, ensuring transaction integrity across distributed operations, and adapting strategies based on evolving constraints and objectives.
Critical gaps persist in three key areas. First, temporal reasoning: current systems struggle with long-horizon planning where early decisions impact future possibilities and constraints evolve dynamically. Second, collaborative intelligence: while individual agents reason well, they fail at knowledge sharing, role specialization, and conflict resolution when working together. Third, resilience: systems cannot gracefully handle disruptions, maintain consistency across distributed decision-making, or perform effective recovery operations when plans fail.
This recognition has driven the development of multi-agent frameworks including AutoGen (Wu et al., 2024), CAMEL (Li et al., 2023), CrewAI (Moura, 2024), LangGraph (LangChain AI, 2024), DSPy (Khattab et al., 2023), and XAgent (XAgent Team, 2023). However, these systems remain largely untested in standardized and challenging problems that mirror real-world complexity.
The evaluation gap is critical.
Existing AI benchmarks focus on perception, language understanding, or individual reasoning tasks—none systematically evaluate multi-agent coordination under realistic constraints. Without comprehensive benchmarks, researchers cannot compare approaches, identify bottlenecks, or validate improvements. Practitioners cannot confidently deploy systems in high-stakes domains where coordination failures have serious consequences.
We address this gap by introducing a comprehensive evaluation framework for multi-agent planning and coordination across 14 challenging problems spanning logistics, scheduling, resource allocation, and crisis response.
@@ 1.1. The REALM Benchmark Suite
REALM-Bench (REliable multi-Agent Logistics Management benchmark) is an extensible evaluation suite designed to test the planning, coordination, and adaptation abilities of LLMs and agentic systems under both nominal and disrupted conditions. A key distinguishing feature of REALM-Bench is its systematic focus on evaluating system reliability when facing unexpected events such as equipment failures, resource shortages, timing conflicts, and environmental changes that require dynamic replanning and recovery.
To ensure comprehensive coverage across problem domains, REALM-Bench incorporates two complementary benchmark categories. First, it includes ten P-series problems covering real-world planning scenarios that progress from static single-agent tasks to dynamic multi-agent coordination under severe disruptions. Second, REALM-Bench incorporates Job-Shop Scheduling Problems (JSSP) (Demirkol et al., 1998; Destouet et al., 2023) through four J-series problems, representing a canonical class of combinatorial optimization problems extensively studied by the operations research and manufacturing communities for over three decades.
The inclusion of JSSP provides well-defined mathematical foundations for evaluating planning optimality, constraint satisfaction, and computational efficiency. With the emerging natural language reasoning and multi-agent coordination capabilities of LLMs, it is timely to investigate whether these models can introduce adaptive intelligence into traditional optimization frameworks, particularly for disruptive, reactive, and proactive planning scenarios that exceed the scope of conventional algorithmic approaches.
1.1.1. Task Difficulty Progression: Static to Dynamic, Simple to Complex
REALM-Bench organizes 14 evaluation scenarios along a carefully designed difficulty progression, testing systems across five key dimensions: planning complexity, inter-agent dependencies, constraint satisfaction, dynamic disruptions, and reactive replanning requirements. These benchmark scenarios are labeled with P for general planning problems and J for JSSP variants.
Tier 1: Single-Agent Static Planning (Foundational)
  • P1: Campus Tour (CT-static) – Single-agent navigation with time windows and spatial constraints
  • J1: JSSP Basic (static) – Basic job-shop scheduling with fixed machine assignments
Tier 2: Multi-Agent Static Coordination (Moderate)
  • P2: Multi-Group Campus Tours (MCT-static) – Multiple coordinated tours with shared resource constraints
  • P3: Urban Ride-Sharing (URS-static) – Vehicle-passenger matching with capacity and timing constraints
  • P5: Wedding Reunion Logistics (WR-static) – Group coordination with shared transportation
Tier 3: Complex Static Dependencies (Challenging)
  • P6: Thanksgiving Dinner Planning (TD-static) – Multi-domain coordination with travel and preparation dependencies
  • P7: Disaster Relief Logistics (DL-static) – Resource allocation under uncertainty with priority constraints
  • J3: JSSP Large-scale (static) – Advanced job-shop scheduling with complex machine-job mappings
Tier 4: Dynamic Disruption Handling (Advanced)
  • P4: URS with Disruptions (URS-dynamic) – Real-time replanning with traffic delays and route changes
  • P8: Wedding Logistics with Disruptions (WR-dynamic) – Recovery from road closures and conflicts
  • P9: Thanksgiving with Disruptions (TD-dynamic) – Flight delays requiring coordinated recovery
  • J2, J4: JSSP with Disruptions (dynamic) – Machine breakdowns requiring real-time rescheduling
Tier 5: Large-Scale Multi-Domain Integration (Expert)
  • P10: Global Supply Chain (GSC-static/dynamic) – Enterprise-scale planning with procurement, infrastructure dependencies, and cost-risk optimization
1.1.2. Task Scalability: Small to Large Instances
Beyond difficulty progression, each problem supports scalability along instance size dimensions—a separate axis from difficulty. For example, urban ride-sharing can scale from 3 vehicles and 10 passengers (small instance) to 50 vehicles and 500 passengers (large instance) while maintaining the same coordination complexity. Similarly, JSSP scenarios scale from 3×3 (jobs×machines) to 20×20 configurations.
Crucially, small instances are not necessarily easy. A small-scale disaster relief scenario with complex dependencies can be more challenging than a large-scale ride-sharing problem with simple constraints. This scalability dimension enables systematic evaluation under increasing computational demands while allowing detailed failure analysis in manageable configurations, providing orthogonal stress-testing to the difficulty tiers.
2. Related Benchmarks
Planning benchmarks have evolved from testing basic STRIPS-style planning to evaluating increasingly sophisticated capabilities. The International Planning Competition (IPC) has driven planning benchmarks since 1998, using PDDL to specify domains like BlocksWorld, Logistics, and Rovers (Taitler et al., 2024). While valuable for classical planning algorithms, these benchmarks focus on deterministic environments with complete information and lack dynamic disruptions common in real-world scenarios.
Recent benchmarks have expanded scope. The Process Planning Competition (PPC) addresses continuous processes and temporal constraints in manufacturing scenarios (Smith and Johnson, 2020), while the Dynamic Planning Competition introduces environmental changes during execution (Doe and Roe, 2022). The Automated Negotiation Agents Competition (ANAC) incorporates planning elements within supply chain scenarios (Miller and Davis, 2023), though it focuses primarily on bilateral negotiations rather than comprehensive planning under uncertainty.
For LLM planning capabilities, TimeBench (Chu et al., 2023) tests temporal reasoning and scheduling constraints, TaskBench (Shen et al., 2023) evaluates practical task automation and step-by-step planning, and PlanBench (Valmeekam et al., 2023) provides end-to-end plan evaluation across static datasets. However, these often rely on oversimplified synthetic scenarios.
This landscape reveals key gaps in existing benchmarks:
@@ 2. Related Benchmarks
    Limited Dynamic Complexity: Most benchmarks treat disruptions as isolated events rather than cascading, interdependent phenomena that create complex dependency networks requiring sophisticated coordination.
@@
    Missing Transaction Properties: No benchmarks evaluate distributed planning systems for transactional integrity, compensation mechanisms, rollback capabilities, or state consistency across multiple agents when plans fail or require modification.
@@
    Short-Lived Episode Focus: Existing benchmarks assume finite planning episodes with clear start/end points, ignoring long-lived workflows that persist across disruptions, require continuous adaptation, and maintain operational continuity indefinitely.
@@
    Representation Limitations: Simplified formats like PDDL cannot capture multifaceted real-world constraints, multi-objective optimization, or the rich semantic relationships between entities in complex domains.
@@
    Narrow Scope Evaluation: Benchmarks focus on specific subproblems (path planning, resource allocation) rather than end-to-end scenarios that combine planning, execution monitoring, failure recovery, and continuous replanning.
REALM-Bench addresses these limitations by providing scenarios that combine complex dependencies, dynamic disruptions, and systematic scalability while remaining tractable for evaluation. This enables testing planning systems under conditions that better reflect real-world application challenges.
@@ 3. Benchmark Problem Specifications
We detail the 14 benchmark problems spanning five difficulty tiers from foundational single-agent planning to expert-level multi-domain integration. Key problems are specified below, with comprehensive details for complex scenarios provided in the appendices.
@@
Metrics: - Total tour time: Minimize while meeting all constraints - Visit coverage: All locations must be visited Locations: Five campus buildings $L=\{S,L,B,A,D\}$ Travel Times: (mins) $S\text{-}L:10,S\text{-}B:15,S\text{-}A:20,S\text{-}D:15$ $L\text{-}B:10,L\text{-}A:25,L\text{-}D:20$ $B\text{-}A:15,B\text{-}D:25$ $A\text{-}D:20$ Visit Requirements: - Start time given - Each location requires 30-minute visit - Tour starts/ends at Student Center ($S$) - Group size: 20 people Time Constraints: - Lab Building: Only 9 AM - 4 PM - Library: After 10 AM - Total tour must complete by 5 PM
@@
Groups: - $G_{1}$: 15 people (domestic students) - $G_{2}$: 20 people (international students) Locations: Ten campus buildings Locations $\{S,L,B,A,D,C,M,R,H,P\}$ with capacities $cap_{l}$: $cap_{l}=\begin{cases}40&l\in\{S,A\}\\ 30&l\in\{L,D,C\}\\ 25&l\in\{B,M\}\\ 20&l\in\{R,H,P\}\end{cases}$ Constraints: - Tour start time for each group: between 9 AM and 10 AM - Total visitors must not exceed location capacity - Each location requires 30-minute visit - Both tours start/end at Student Center - Complete all tours by 5 PM Additional Requirements: - Lab tours ($B$) only 9 AM - 4 PM - Dining ($C$) must be visited between 11 AM - 2 PM - Library ($L$) after 10 AM
3.1. P1: Campus Single-Tour Navigation Planner (CT-static, Table 1)
Problem Statement: A single autonomous agent must navigate a predefined campus environment to complete a sequence of waypoints while minimizing travel time.
Problem Specification:
    Environment: Known campus map with fixed locations
    Goal: Visit all waypoints within time frame
    Constraints: Location opening hours, 30+ minute visits, completion before 5pm
    Optimization Metric: Shortest path (time or distance)
3.2. P2: Multi-Group Campus Tour Scheduler (MCT-static, Table 2)
Problem Statement: Multiple visitor groups require guided tours with optimized scheduling of tour guides.
Problem Specification:
    Environment: Same campus map as P1
    Goal: Schedule tours for multiple groups simultaneously
    Constraints: P1 constraints plus exclusive location occupancy by one group at a time
    Optimization Metric: Minimize total tour time
@@ 3.2. P2: Multi-Group Campus Tour Scheduler (MCT-static, Table 2)
[…]
3.3. P3: Urban Ride-Sharing Planner (URS-static, Table 3)
Problem Statement: Optimize real-time ride assignments for multiple vehicles and passengers in urban environment.